### v7

@@ 3.3 Transferablity analysis
Besides the positive confirmation above, in this section, we provide quantitative corroboration for our claim that top layers might not be needed in TL for MIC. To this end, we need a variant of the classical canonical correlation analysis (CCA), singular-vector CCA (SVCCA) (Raghu et al.(2017)Raghu, Gilmer, Yosinski, and Sohl-Dickstein).
@@
which computes the first $k$ pairs of most correlated linear combinations. In practice, all the covariance matrices $\mathbf{\Sigma}_{\mathbf{x}\mathbf{y}}$, $\mathbf{\Sigma}_{\mathbf{x}\mathbf{x}}$, and $\mathbf{\Sigma}_{\mathbf{y}\mathbf{y}}$ are replaced by their finite-sample approximations, and the covariances of the top $k$ most correlated pairs are the top $k$ singular values of $\mathbf{\Sigma}^{-1/2}_{\mathbf{x}\mathbf{x}}\mathbf{\Sigma}_{\mathbf{x}\mathbf{y}}\mathbf{\Sigma}^{-1/2}_{\mathbf{y}\mathbf{y}}$ (Uurtio et al.(2018)Uurtio, Monteiro, Kandola, Shawe-Taylor, Fernandez-Reyes, and Rousu), all of which lie in $[0,1]$. We call these singular values the CCA coefficients. Obviously, high values in these coefficients indicate high levels of correlation.
@@
Features in top layers become more correlated with bottom layers after TL The “horizontal” analysis above says the features of the top layers are almost re-learned in FTL, LWFT, and TF, but it remains unclear what features are learned there. If we believe that high-level features are probably not useful for COVID classification (Shi et al.(2020)Shi, Han, Jiang, Cao, Alwalid, Gu, Fan, and Zheng), a reasonable hypothesis is that these top layers actually learn features that are more correlated with those of lower layers after TL. This seems indeed the case, as shown in Fig. 4 (ii): the correlation level between the block2 and block 4 features, as well as between block3 and block4 features, increases both visibly and quantitatively.
@@ 4 Experiments
Experiment setup We systematically compare FTL, TF, LWFT, and our TTL on $3$ MIC tasks covering both 2D and 3D image modalities, and also explore a 2D lung segmentation task. For the 3D MIC task, we choose ResNeXt3D-101 as the default model and compare it with PENet which is a handcrafted model in Huang et al.(2020a)Huang, Kothari, Banerjee, Chute, Ball, Borus, Huang, Patel, Rajpurkar, Irvin, Dunnmon, Bledsoe, Shpanskaya, Dhaliwal, Zamanian, Ng, and Lungren. Both models are first pretrained on the kinetics-600 dataset (Carreira et al.(2018)Carreira, Noland, Banki-Horvath, Hillier, and Zisserman), and then finetuned with the same setting as in Huang et al.(2020a)Huang, Kothari, Banerjee, Chute, Ball, Borus, Huang, Patel, Rajpurkar, Irvin, Dunnmon, Bledsoe, Shpanskaya, Dhaliwal, Zamanian, Ng, and Lungren: $0.01$ initial LR for randomly initialized weights and $0.1$ for pretrained weights, SGD (momentum $=0.9$) optimizer, cosine annealing LR scheduler, $100$ epochs of training, and best model selected based on the validation AUROC. For experiments involving randomness, we repeat them $3$ times and report the mean and standard deviation. More detailed experiment setup, ablation studies, and other explorations can be found in supplementary materials.
@@
COVID-19 chest x-ray image classification We take the chest x-rays from the BIMCV-COVID19$+$ (containing COVID positives) and BIMCV-COVID19-- (containing COVID negatives) datasets (iteration 1) (Vayá et al.(2020)Vayá, Saborit, Montell, Pertusa, Bustos, Cazorla, Galant, Barber, Orozco-Beltrán, García-García, et al.) [^4] https://bimcv.cipf.es/bimcv-projects/bimcv-covid19/#1590858128006-9e640421-6711.. We manually remove the small number of lateral views and outliers, leaving $2261$ positives and $2463$ negatives for our experiment. We have demonstrated the plausibility and superiority of TTL based on this MIC task around Section 3.2, which we do not repeat here.
Mitotic cells classification The density of mitotic cells undergoing division (i.e., mitotic figures) is known to be related to tumor proliferation and can be used for tumor prognosis (Aubreville et al.(2022)Aubreville, Bertram, Veta, Breininger, Jabari, and Stathonikos). Since cell division changes its morphology, we expect mid-level blob-like features to be the determinant here. We take the dataset from the mitotic domain generalization challenge (MIDOG2022) (Aubreville et al.(2022)Aubreville, Bertram, Veta, Breininger, Jabari, and Stathonikos) which is about detection of mitotic figures, i.e., the training set consists of properly cropped $9501$ mitotic figures and $11051$ non-mitotic figures, and the task is to localize mitotic figures on large pathological slices during the test. We modify the task into a binary MIC (positives are mitotic figures, and negatives are non-mitotic figures).
@@
Pulmonary embolism CT image classification Pulmonary Embolism (PE) is a blockage of the blood vessels connecting the lungs and the heart, and CT pulmonary angiography (CTPA) is the gold standard for its diagnosis (Wittram et al.(2004)Wittram, Maher, Yoo, Kalra, Shepard, and McLoud). We take the public PE dataset (Huang et al.(2020a)Huang, Kothari, Banerjee, Chute, Ball, Borus, Huang, Patel, Rajpurkar, Irvin, Dunnmon, Bledsoe, Shpanskaya, Dhaliwal, Zamanian, Ng, and Lungren) consisting of $1797$ CT images from $1773$ patients, and compare the performance with handcrafted 3DCNN model, PENet which is the SOTA model that outperforms other 3D DCNN models, such as ResNet3D-50, ResNeXt3D-101, and DenseNet3D-121 as demonstrated in Huang et al.(2020a)Huang, Kothari, Banerjee, Chute, Ball, Borus, Huang, Patel, Rajpurkar, Irvin, Dunnmon, Bledsoe, Shpanskaya, Dhaliwal, Zamanian, Ng, and Lungren.
On CT images, PE often appears as localized blobs that map to mid-level visual features. So the suboptimal TL performance of the SOTA models is mostly likely due to the rigid FTL strategy. We confirm this on ResNeXt3D-101 pretrained on kinetics-600: after layer-wise search, all differential TL methods including TF, LWFT, and TTL outperform PENet by considerable margins in both AUROC and AUPRC, as shown in Table 2 (bottom left). Also, both TTL-1 and TF-1 find the best cutoff at the transition of block3 and block4, another confirmation of our intuition that probably only low- to mid-level features are needed here. We note that although the AUROC obtained via FTL is slightly lower than that of PENet, the AUPRC is actually higher—which (Huang et al.(2020a)Huang, Kothari, Banerjee, Chute, Ball, Borus, Huang, Patel, Rajpurkar, Irvin, Dunnmon, Bledsoe, Shpanskaya, Dhaliwal, Zamanian, Ng, and Lungren) does not consider when drawing their conclusion. Our TTL-2 is a clear winner in performance, despite that PENet has been meticulously designed and optimized for the task.
Chest X-ray lung segmentation Despite our focus on MIC, we briefly explore the potential of TTL for segmentation also. To this end, we explore a public chest x-ray lung segmentation datasets collected from two source: Montgomery Country XCR set (MC) and Shenzhen Hospital CXR Set (SH) (Jaeger et al.(2013)Jaeger, Karargyris, Candemir, Folio, Siegelman, Callaghan, Xue, Palaniappan, Singh, Antani, et al.; Candemir et al.(2013)Candemir, Jaeger, Palaniappan, Musco, Singh, Xue, Karargyris, Antani, Thoma, and McDonald), both of which provide manual segmentation masks. MC consists of $58$/$80$ tuberculosis/normal cases, and SH has $336$/$326$ tuberculosis/normal cases.
@@ A.2 Extending TTL on segmentation models
Figure 5: Similar to TTL on ResNet/DenseNet, we identify the block structure in U-Net and truncate at the intersection between blocks. The image of U-Net is adapted from Ronneberger et al.(2015)Ronneberger, Fischer, and Brox
@@ B.1 Impact of network architecture
[…]
@@ C.2 Inter-domain TL
All we have discussed above are intra-domain TL where the source domain is distinct from the target domain. It is tempting to think inter-domain TL might be more effective. To test this, we take the x-ray-based COVID-19 classification task again, but now finetune the ResNet50 model pretrained on CheXpert, a massive-scale x-ray dataset ($\sim 220K$ images) for disease prediction covering $13$ types of diseases (see Table 4(Irvin et al.(2019)Irvin, Rajpurkar, Ko, Yu, Ciurea-Ilcus, Chute, Marklund, Haghgoo, Ball, Shpanskaya, et al.). These diseases correspond to varying levels of visual features. Hence, we categorize them into two groups: CheXpert-low includes diseases that need low-level features only, and CheXpert-high covers those needing both low- and high-level features. A summary of the categorization can be found in Table 4555Disease symptoms information obtained from Mayo clinic: https://www.mayoclinic.org/symptom-checker/select-symptom/itt-20009075. We pretrain ResNet50 on $3$ variants of the dataset, respectively: full CheXpert, CheXpert-low, and CheXpert-high, and then compare FTL and our TTL on the three resulting models. For TTL, we only perform a coarse-scale search and pick the $3$ transition layers between the $4$ blocks in ResNet50 as truncation points. As always, all experiments are repeated three times.
@@ Appendix D Sample images of Grad-CAM analysis
To find out what visual features are learned by different models, we conduct a feature analysis study using Grad-CAM(Selvaraju et al.(2017)Selvaraju, Cogswell, Das, Vedantam, Parikh, and Batra). Our empirical findings show that deep models tend to produce spatially extensive features that spread over the image. In contrast, shallow models such as those produced by TTL tend to find spatially localized features that often capture the infected area around the lungs and the mitotic region, as shown in Fig. 1 and Table 2. In particular, those localized regions often take the form of the abnormal texture of boundary changes and are invariant to the global object (e.g., the shape of the ribs and cells), which demonstrates the effectiveness of TTL in learning more discriminative features for MIC.

### v8

@@ 3.3 Transferablity analysis
Besides the positive confirmation above, in this section, we provide quantitative corroboration for our claim that top layers might not be needed in TL for MIC. To this end, we need a variant of the classical canonical correlation analysis (CCA), singular-vector CCA (SVCCA) (Raghu et al., 2017).
@@
which computes the first $k$ pairs of most correlated linear combinations. In practice, all the covariance matrices $\mathbf{\Sigma}_{\mathbf{x}\mathbf{y}}$, $\mathbf{\Sigma}_{\mathbf{x}\mathbf{x}}$, and $\mathbf{\Sigma}_{\mathbf{y}\mathbf{y}}$ are replaced by their finite-sample approximations, and the covariances of the top $k$ most correlated pairs are the top $k$ singular values of $\mathbf{\Sigma}^{-1/2}_{\mathbf{x}\mathbf{x}}\mathbf{\Sigma}_{\mathbf{x}\mathbf{y}}\mathbf{\Sigma}^{-1/2}_{\mathbf{y}\mathbf{y}}$ (Uurtio et al., 2018), all of which lie in $[0,1]$. We call these singular values the CCA coefficients. Obviously, high values in these coefficients indicate high levels of correlation.
@@
Features in top layers become more correlated with bottom layers after TL The “horizontal” analysis above says the features of the top layers are almost re-learned in FTL, LWFT, and TF, but it remains unclear what features are learned there. If we believe that high-level features are probably not useful for COVID classification (Shi et al., 2020), a reasonable hypothesis is that these top layers actually learn features that are more correlated with those of lower layers after TL. This seems indeed the case, as shown in Fig. 4 (ii): the correlation level between the block2 and block 4 features, as well as between block3 and block4 features, increases both visibly and quantitatively.
@@ 4 Experiments
Experiment setup We systematically compare FTL, TF, LWFT, and our TTL on $3$ MIC tasks covering both 2D and 3D image modalities, and also explore a 2D lung segmentation task. For the 3D MIC task, we choose ResNeXt3D-101 as the default model and compare it with PENet which is a handcrafted model in Huang et al. (2020a). Both models are first pretrained on the kinetics-600 dataset (Carreira et al., 2018), and then finetuned with the same setting as in Huang et al. (2020a): $0.01$ initial LR for randomly initialized weights and $0.1$ for pretrained weights, SGD (momentum $=0.9$) optimizer, cosine annealing LR scheduler, $100$ epochs of training, and best model selected based on the validation AUROC. For experiments involving randomness, we repeat them $3$ times and report the mean and standard deviation. More detailed experiment setup, ablation studies, and other explorations can be found in supplementary materials.
@@
COVID-19 chest x-ray image classification We take the chest x-rays from the BIMCV-COVID19$+$ (containing COVID positives) and BIMCV-COVID19-- (containing COVID negatives) datasets (iteration 1) (Vayá et al., 2020) [^4] https://bimcv.cipf.es/bimcv-projects/bimcv-covid19/#1590858128006-9e640421-6711.. We manually remove the small number of lateral views and outliers, leaving $2261$ positives and $2463$ negatives for our experiment. We have demonstrated the plausibility and superiority of TTL based on this MIC task around Section 3.2, which we do not repeat here.
Mitotic cells classification The density of mitotic cells undergoing division (i.e., mitotic figures) is known to be related to tumor proliferation and can be used for tumor prognosis (Aubreville et al., 2022). Since cell division changes its morphology, we expect mid-level blob-like features to be the determinant here. We take the dataset from the mitotic domain generalization challenge (MIDOG2022) (Aubreville et al., 2022) which is about detection of mitotic figures, i.e., the training set consists of properly cropped $9501$ mitotic figures and $11051$ non-mitotic figures, and the task is to localize mitotic figures on large pathological slices during the test. We modify the task into a binary MIC (positives are mitotic figures, and negatives are non-mitotic figures).
@@
Pulmonary embolism CT image classification Pulmonary Embolism (PE) is a blockage of the blood vessels connecting the lungs and the heart, and CT pulmonary angiography (CTPA) is the gold standard for its diagnosis (Wittram et al., 2004). We take the public PE dataset (Huang et al., 2020a) consisting of $1797$ CT images from $1773$ patients, and compare the performance with handcrafted 3DCNN model, PENet which is the SOTA model that outperforms other 3D DCNN models, such as ResNet3D-50, ResNeXt3D-101, and DenseNet3D-121 as demonstrated in Huang et al. (2020a).
On CT images, PE often appears as localized blobs that map to mid-level visual features. So the suboptimal TL performance of the SOTA models is mostly likely due to the rigid FTL strategy. We confirm this on ResNeXt3D-101 pretrained on kinetics-600: after layer-wise search, all differential TL methods including TF, LWFT, and TTL outperform PENet by considerable margins in both AUROC and AUPRC, as shown in Table 2 (bottom left). Also, both TTL-1 and TF-1 find the best cutoff at the transition of block3 and block4, another confirmation of our intuition that probably only low- to mid-level features are needed here. We note that although the AUROC obtained via FTL is slightly lower than that of PENet, the AUPRC is actually higher—which (Huang et al., 2020a) does not consider when drawing their conclusion. Our TTL-2 is a clear winner in performance, despite that PENet has been meticulously designed and optimized for the task.
Chest X-ray lung segmentation Despite our focus on MIC, we briefly explore the potential of TTL for segmentation also. To this end, we explore a public chest x-ray lung segmentation datasets collected from two source: Montgomery Country XCR set (MC) and Shenzhen Hospital CXR Set (SH) (Jaeger et al., 2013; Candemir et al., 2013), both of which provide manual segmentation masks. MC consists of $58$/$80$ tuberculosis/normal cases, and SH has $336$/$326$ tuberculosis/normal cases.
@@ A.2 Extending TTL on segmentation models
Figure 5: Similar to TTL on ResNet/DenseNet, we identify the block structure in U-Net and truncate at the intersection between blocks. The image of U-Net is adapted from Ronneberger et al. (2015)
@@ B.1 Impact of network architecture
  Method AUROC AUPRC Params(M) MACs(G) Speed(s) CPU GPU RN501 FTL $0.853\pm 0.004$ $0.861\pm 0.002$ 23.5 4.12 80.0 6.72 TTL $\mathbf{0.865\pm 0.001}$ $\mathbf{0.870\pm 0.007}$ 8.55 3.31 60.9 5.83 DN1212 FTL $0.852\pm 0.002$ $0.856\pm 0.004$ 18.2 3.31 107 24.0 TTL $\mathbf{0.866\pm 0.006}$ $\mathbf{0.871\pm 0.012}$ 1.53 2.11 46.2 5.94 ENb03 TL $0.795\pm 0.004$ $0.794\pm 0.002$ 3.63 0.378 30.0 8.33 TTL $\mathbf{0.842\pm 0.001}$ $\mathbf{0.843\pm 0.007}$ 0.896 0.255 24.0 6.46   Abbreviation: 1 ResNet50, 2:DenseNet121, 3 EfficientNet B0
@@ C.2 Inter-domain TL
All we have discussed above are intra-domain TL where the source domain is distinct from the target domain. It is tempting to think inter-domain TL might be more effective. To test this, we take the x-ray-based COVID-19 classification task again, but now finetune the ResNet50 model pretrained on CheXpert, a massive-scale x-ray dataset ($\sim 220K$ images) for disease prediction covering $13$ types of diseases (see Table 4(Irvin et al., 2019). These diseases correspond to varying levels of visual features. Hence, we categorize them into two groups: CheXpert-low includes diseases that need low-level features only, and CheXpert-high covers those needing both low- and high-level features. A summary of the categorization can be found in Table 4555Disease symptoms information obtained from Mayo clinic: https://www.mayoclinic.org/symptom-checker/select-symptom/itt-20009075. We pretrain ResNet50 on $3$ variants of the dataset, respectively: full CheXpert, CheXpert-low, and CheXpert-high, and then compare FTL and our TTL on the three resulting models. For TTL, we only perform a coarse-scale search and pick the $3$ transition layers between the $4$ blocks in ResNet50 as truncation points. As always, all experiments are repeated three times.
@@ Appendix D Sample images of Grad-CAM analysis
To find out what visual features are learned by different models, we conduct a feature analysis study using Grad-CAM(Selvaraju et al., 2017). Our empirical findings show that deep models tend to produce spatially extensive features that spread over the image. In contrast, shallow models such as those produced by TTL tend to find spatially localized features that often capture the infected area around the lungs and the mitotic region, as shown in Fig. 1 and Table 2. In particular, those localized regions often take the form of the abnormal texture of boundary changes and are invariant to the global object (e.g., the shape of the ribs and cells), which demonstrates the effectiveness of TTL in learning more discriminative features for MIC.